# SWE-Dev: Evaluating and Training Autonomous Feature-Driven Software Development

### v2

@@ C.1 Analysis of PRD Quality
C.2 Explanation on the Underperformance of Reasoning Models
@@
C.3 Error Analysis
@@
C.4 Limitation and Future Work
@@
Our training experiments focus on standard techniques—SFT, RL, and role-wise MAS training—which yield modest gains. Future work could explore stronger RL [45, 46, 47], dynamic agent coordination [48], and curriculum learning [49]. Notably, SWE-Dev offers fine-grained complexity signals via call trees that can guide complexity-aware training.
C.5 Broader impacts
@@
Appendix D Detailed Benchmark Construction
D.1 Call tree generation
@@
D.2 Docstring Augmentation Prompt
@@
Appendix E Extended Related Work
E.1 Multi-agent system
@@
Appendix F Experiment Settings
F.1 Inference
LLMs. We evaluate 17 chatbot LLMs with different model size, including Qwen2.5-Instruct models 1.5B/3B/7B/14B/32B/72B [50], Qwen3 models 8B/30B-A3B [51], Llama 3.1-8B/3.3-70B-Instruct [52], Phi 4 [53], Claude-3.7-Sonnet [11], Deepseek-V3 [54], GPT-4o [12], Deepseek-Coder-V2-Lite-Instruct [55], Qwen2.5-Coder-14B-Instruct [56]. Additionally, We extend the evaluation to reasoning models, including Deepseek-R1-distill models (Qwen2.5 7B/32B, Llama-70B) [57], Qwen3 8B/30B-A3B (thinking) [51], QwQ-32B-Preview, Deepseek-R1 [57], OpenAI-o1 [58], Claude-3.7-Sonnet-thinking [11], and Grok-3-Beta [59].
Multi-Agent Systems. To provide a more comprehensive evaluation of SWE-Dev, we expand our study to include multi-agent systems (MAS) built on LLMs. Prior research has demonstrated that MAS can enhance performance on tasks requiring multi-step reasoning and coordination [48, 60, 40]. In our experiments, all MAS are implemented using GPT-4o-mini [61] as the underlying model to ensure consistency across methods. And for fair comparison, we utilize MASLab [62], a unified framework integrating multiple MAS implementations. We evaluate coordination-based MAS such as LLM Debate [36], Self Refine [34], Multi-Agent Debate (MAD) [37], and Self Consistency [35] that feature relatively simple agent interaction strategies. We further include structured, workflow-oriented MAS designed for code generation, including Agentverse [38], MetaGPT [60], ChatDev [40], MapCoder [39], and EvoMAC [18].
F.2 Training
F.3 Single-Agent Supervied Fine-tuning
@@
F.4 Single-Agent Reinforcement Learning
For reinforcement learning (RL) training, we sampled 2k instances from the SWE-Dev to balance computational feasibility and the ability to capture RL benefits. Specifically, we used Proximal Policy Optimization (PPO) [42] and Direct Preference Optimization (DPO) [43] for training the Qwen2.5-7B-Instruct using 8 NVIDIA A100 GPUs.
PPO: The training was conducted using a batch size of 256 and trained for 5 epochs, with a learning rate of $1\times 10^{-6}$ for the actor and $1\times 10^{-5}$ for the critic. The training was set to save checkpoints every 10 steps. We used a maximum prompt length of 8192 tokens and set a micro-batch size of 32. The reward for PPO is calculated based on the pass rate of the test cases.
@@ F.4 Single-Agent Reinforcement Learning
For a fair comparison with SFT in §4.2.2, we used the same 2k training samples for both SFT and RL. The details for SFT training are outlined in Appendix F.3.
@@
F.5 Mingle-Agent Supervied Fine-tuning
In our multi-agent fine-tuning experiments, we utilize a simplified version of EvoMAC [18], retaining only two core roles: Organizer and Coder, see Figure 17. The fine-tuning process follows an iterative workflow. Initially, the Organizer processes the Project Requirement Description (PRD) and breaks it down into clearly defined subtasks or instructions. Subsequently, the Coder generates corresponding code implementations for these subtasks. The generated code is then evaluated using the provided ground truth (GT) test cases. Feedback from these evaluations informs subsequent iterations, enabling iterative refinement of both the task decomposition by the Organizer and the code generation by the Coder. Fig. 19 and Fig. 20 respectively show the prompts used for the Coder and Organizer.
@@ F.5 Mingle-Agent Supervied Fine-tuning
To effectively leverage the feedback from GT test cases, we employ rejection sampling—a method widely adopted in reinforcement learning and language model fine-tuning [63, 64]. The detailed procedure is as follows:

### v3

@@ E.2 Analysis of PRD Quality
E.3 Analysis on the Underperformance of Reasoning Models
@@
E.4 Analysis of MAS
While multi-agent systems (MAS) demonstrate performance gains over single-agent baselines on SWE-Dev, their behaviors vary significantly across designs. To better understand these differences, we analyze MAS from a communication and coordination perspective. Specifically, we examine how agent structure and interaction overhead relate to execution success on repository-level feature development tasks.
MAS System Pass@1 (%) Avg. LLM Calls # Roles Avg. Agent-to-Agent Msgs
Reflexion 39.77 2.12 3 10.77
Self-Refine 40.02 5.00 2 26.79
Self-Consistency 37.62 6.00 2 33.01
LLM Debate 38.48 7.00 2 76.35
MAD 31.50 7.00 4 2.36
AgentVerse 38.67 4.52 4 8.83
MetaGPT 29.56 9.69 6 24.26
MapCoder 24.55 21.01 5 63.03
Table 11: Communication and coordination statistics of MAS on SWE-Dev. All systems are driven by the base model (GPT-4o-mini). We report Pass@1, average number of LLM calls per task, number of predefined agent roles, and the average number of agent-to-agent messages.
Table 11 reveals two consistent trends. First, simple and low-overhead MAS achieve the best performance–efficiency trade-off. General-purpose methods such as Reflexion [reflexion] and Self-Refine [self_refine] use only two to three roles and require relatively few model calls. Despite their simplicity, they reach the highest Pass@1 scores. For example, Reflexion achieves 39.77% Pass@1 with only 2.12 LLM calls and 10.77 agent-to-agent messages on average. These systems rely on iterative self-refinement rather than explicit division of labor. As a result, the global task context remains concentrated within a small number of agents, reducing coordination cost.
Second, workflow-heavy and code-specific MAS suffer from coordination overhead and context fragmentation. Systems such as MetaGPT [hong2024metagptmetaprogrammingmultiagent] and MapCoder [islam2024mapcodermultiagentcodegeneration] introduce more roles, more calls, and substantially more inter-agent communication. However, this increased interaction does not translate into better execution success. MapCoder, for instance, uses over 21 LLM calls and exchanges more than 60 agent-to-agent messages on average, yet achieves the lowest Pass@1 among all compared methods. Each additional handoff between agents introduces opportunities for partial context loss and error propagation. This effect is particularly harmful on SWE-Dev tasks, which already involve long code contexts and multi-file dependencies.
Overall, these results suggest that, for repository-level feature development, maintaining a stable and unified reasoning context is more important than fine-grained role specialization. Simple self-refinement style MAS exhibit an emergent robustness by minimizing communication overhead while preserving global task coherence. In contrast, heavily engineered multi-role workflows tend to accumulate coordination cost without corresponding performance gains.
E.5 Error Analysis
@@
E.6 Limitation and Future Work
@@
Our training experiments focus on standard techniques—SFT, RL, and role-wise MAS training—which yield modest gains. Future work could explore stronger RL [grpo, dapo, reinforce++], dynamic agent coordination [masgpt], and curriculum learning [bengio2009curriculum]. Notably, SWE-Dev offers fine-grained complexity signals via call trees that can guide complexity-aware training.
E.7 Broader impacts
@@
Appendix F Detailed Benchmark Construction
F.1 Call tree generation
@@
F.2 Docstring Augmentation Prompt
@@
F.3 Masking Procedure and Reproducibility
Our dataset construction does not rely on any task-complexity thresholds such as call-tree depth or node count. Instead, we apply a fixed set of masking rules uniformly across all repositories. Specifically, for each task we derive the dynamic call chain exercised by the upstream test and construct a masked input using one of three modes: (i) masking depth-1 target functions, (ii) masking depth-2 dependent functions, or (iii) masking the entire call chain for call-tree depth $\geq$ 2. These masking modes are predetermined and do not involve any further filtering or parameter tuning.
Appendix G Extended Related Work
G.1 Multi-agent system
@@
Appendix H Experiment Settings
H.1 Inference
LLMs. We evaluate 17 chatbot LLMs with different model size, including Qwen2.5-Instruct models 1.5B/3B/7B/14B/32B/72B [qwen2025qwen25technicalreport], Qwen3 models 8B/30B-A3B [qwen3], Llama 3.1-8B/3.3-70B-Instruct [grattafiori2024llama3herdmodels], Phi 4 [abdin2024phi4technicalreport], Claude-3.7-Sonnet [claude3.7], Deepseek-V3 [deepseekai2025deepseekv3technicalreport], GPT-4o [gpt4o], Deepseek-Coder-V2-Lite-Instruct [deepseekai2024deepseekcoderv2breakingbarrierclosedsource], Qwen2.5-Coder-14B-Instruct [hui2024qwen25codertechnicalreport]. Additionally, We extend the evaluation to reasoning models, including Deepseek-R1-distill models (Qwen2.5 7B/32B, Llama-70B) [deepseek], Qwen3 8B/30B-A3B (thinking) [qwen3], QwQ-32B-Preview, Deepseek-R1 [deepseek], OpenAI-o1 [openai2024openaio1card], Claude-3.7-Sonnet-thinking [claude3.7], and Grok-3-Beta [grok3].
Multi-Agent Systems. To provide a more comprehensive evaluation of SWE-Dev, we expand our study to include multi-agent systems (MAS) built on LLMs. Prior research has demonstrated that MAS can enhance performance on tasks requiring multi-step reasoning and coordination [masgpt, hong2024metagptmetaprogrammingmultiagent, qian2024chatdevcommunicativeagentssoftware]. In our experiments, all MAS are implemented using GPT-4o-mini [gpt4omini] as the underlying model to ensure consistency across methods. And for fair comparison, we utilize MASLab [ye2025maslabunifiedcomprehensivecodebase], a unified framework integrating multiple MAS implementations. We evaluate coordination-based MAS such as LLM Debate [llm_debate], Self Refine [self_refine], Multi-Agent Debate (MAD) [mad], and Self Consistency [self_consistency] that feature relatively simple agent interaction strategies. We further include structured, workflow-oriented MAS designed for code generation, including Agentverse [chen2023agentversefacilitatingmultiagentcollaboration], MetaGPT [hong2024metagptmetaprogrammingmultiagent], ChatDev [qian2024chatdevcommunicativeagentssoftware], MapCoder [islam2024mapcodermultiagentcodegeneration], and EvoMAC [evomac].
H.2 Training
H.3 Single-Agent Supervised Fine-tuning
@@
H.4 Single-Agent Reinforcement Learning
[…]
PPO: The training was conducted using a batch size of 256 and trained for 5 epochs, with a learning rate of $1\times 10^{-6}$ for the actor and $1\times 10^{-5}$ for the critic. We set the KL coefficient to 0.001. The training was set to save checkpoints every 10 steps. We used a maximum prompt length of 8192 tokens and set a micro-batch size of 32. The maximum response length is 2048 tokens. The reward for PPO is calculated based on the pass rate of the test cases. Concretely, for each rollout solution, the reward is defined as the pass rate, i.e., the number of test cases successfully passed by the solution divided by the total number of test cases available for that task. During RL training, we do not employ any external agent framework: the language model directly receives the problem requirement description (PRD) together with the relevant code context as input and generates a single rollout solution.
@@ H.4 Single-Agent Reinforcement Learning
For a fair comparison with SFT in §4.2.2, we used the same 2k training samples for both SFT and RL. The details for SFT training are outlined in Appendix H.3.
@@
Figure 16: The PPO training progress of Qwen2.5-7B-Instrut on the SWE-Dev with 40 steps and 2048 max token length. The light curve shows the raw reward, while the dark curve applies EMA smoothing.
Figure 17: The PPO training progress of Qwen2.5-7B-Instrut on the SWE-Dev with 160 steps and 8192 max token length. The light curve shows the raw reward, while the dark curve applies EMA smoothing.
Figure 18: Entropy trajectory during PPO training (40 updates). The light curve shows the raw policy entropy, and the dark curve shows the EMA-smoothed trend. Entropy decreases steadily over training, indicating reduced exploration and a progressively sharper output distribution—consistent with the observed contraction in diversity under RL.
H.5 Multi-Agent Supervised Fine-tuning
In our multi-agent fine-tuning experiments, we utilize a simplified version of EvoMAC [evomac], retaining only two core roles: Organizer and Coder, see Figure 19. The fine-tuning process follows an iterative workflow. Initially, the Organizer processes the Project Requirement Description (PRD) and breaks it down into clearly defined subtasks or instructions. Subsequently, the Coder generates corresponding code implementations for these subtasks. The generated code is then evaluated using the provided ground truth (GT) test cases. Feedback from these evaluations informs subsequent iterations, enabling iterative refinement of both the task decomposition by the Organizer and the code generation by the Coder. Fig. 21 and Fig. 22 respectively show the prompts used for the Coder and Organizer.
@@ H.5 Multi-Agent Supervised Fine-tuning
To effectively leverage the feedback from GT test cases, we employ rejection sampling—a method widely adopted in reinforcement learning and language model fine-tuning [zhou2024archertraininglanguagemodel, snell2023offlinerlnaturallanguage]. The detailed procedure is as follows: